OpenStudy (anonymous):

find an equation of the tangent line to the graph of y = g(x) at x=5 if g(5) = -3 and g`(5) = 4

Miracrown (miracrown):

g'(5)=4, is that what you wrote?

Miracrown (miracrown):

The slope of the line is g'(5)=4. (5,-3) is the point on the line. So the line is y+3=5(x-5) or y=5x-28.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

wait, wouldn't it be y+3=4(x-5) or am I getting myself confused on this?